40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 10068     Entry time: Thu Jun 19 00:02:48 2014
Author: Jenne 
Type: Update 
Category: LSC 
Subject: PRMI+2 arms recovery 

Arms locked in comm and diff with ALS. PRMI locked with REFL33 I&Q while arms off resonance. Having trouble reducing CARM offset, even to get to arm powers of 1.

After Manasa installed the new Xgreen PD, Koji looked at the PSL table alignment with me.  I saw only a very weak beatnote with the X BBPD, even though I could see the beatnote on the Y PD from the leakage of the X beam to the Y PD  (Yend shutter was closed, so just PSL and X greens were on the table).  I had thought that my near-field and far-field alignments were pretty good (actually, I checked them, but didn't feel that I needed to tweak them since Manasa did the alignment this afternoon).  Anyhow, it was just a matter of tweaking up the alignment a bit, and then the X beatnote got up to about -25dBm at a few tens of MHz.  I am starting to question myself if the other BBPD is broken, or if I just not well enough aligned.  Anyhow, the spare is in, we can still have a look at the previous X BBPD, but it may be okay, and it's just me embarrassing myself by not catching an alignment problem.

Anyhow, after the X beat was found, I was able to (on my first try) lock the arms using ALS comm and diff.  (I already had a nice strong Y beatnote, so that didn't need finding, other than temp adjustment of the end laser).  I ran the carm_cm_up.sh script, and it did everything nicely.  I did a quickie check of the phase tracker loop gains, but that should be redone in the morning.

PRMI was a little reluctant to lock, so I played around with the MICH and PRCL gains, but didn't really find any combination that was any better than the usual (+0.8 for MICH, -0.04 for PRCL as I had last night, although I needed to reduce the PRCL gain back to -0.02 to eliminate loop osc). 

After an arm lockloss, I relocked just the PRMI and used awggui to put a line into C1:LSC-PRM_EXC to check the RF PD phasing.  I changed REFL33 from 133.5 to 138.5, and REFL 165 from -142.5 to -152.5.  I didn't think that REFL11 needed changing, and I didn't check REFL55.  I also checked that I could lock PRMI without arms, using both REFL33 and REFL165 - they seemed about the same to me, both stable.  REFL33 has 1's in the input matrix, and I was using 0.07's for REFL165.  The demod phase adjustment didn't really improve PRMI locking while the arms were held off resonance, even if I moved the arms even farther from resonance (usually we do 3nm, I went out to 5nm to see if that helped - it didn't).  I tried REFL165 locking, but that wasn't any good either.  I tried using REFL165 with the arms held off resonance, but that didn't seem to catch at all (at least with REFL33 I was getting short lock blips). 

Anyhow, of the 3 or 4 times that I caught REFL33 PRMI lock and tried to reduce the CARM offset, only one time did I even get to arm powers of about 1 (CARM digital offset of -0.1, with CARM held on sqrt transmission signals), and then it didn't stay for more than a few tens of seconds.  The other few times, it broke lock on the way up to arm powers of 1. 

So, carm_cm_up.sh works pretty well, although perhaps the arm powers of 1 offset reduction needs to be a little slower.  PRMI doesn't catch and hold lock very easily with REFL33, and even less so with REFL165.  It may be useful to try catching lock with REFL11 or 55, and doing a transition over to 3f.  No real progress forward, but we're pretty much recovered.

 

ELOG V3.1.3-